# DISCOVER: Automated Curricula for Sparse-Reward Reinforcement Learning

### v1

@@ Abstract
Sparse-reward reinforcement learning (RL) can model a wide range of highly complex tasks. Solving sparse-reward tasks is RL’s core premise — requiring efficient exploration coupled with long-horizon credit assignment — and overcoming these challenges is key for building self-improving agents with superhuman ability. We argue that solving complex and high-dimensional tasks requires solving simpler tasks that are relevant to the target task. In contrast, most prior work designs strategies for selecting exploratory tasks with the objective of solving any task, making exploration of challenging high-dimensional, long-horizon tasks intractable. We find that the sense of direction, necessary for effective exploration, can be extracted from existing RL algorithms, without needing any prior information. Based on this finding, we propose a method for directed sparse-reward goal-conditioned very long-horizon RL (DISCOVER), which selects exploratory goals in the direction of the target task. We connect DISCOVER to principled exploration in bandits, formally bounding the time until the target task becomes achievable in terms of the agent’s initial distance to the target, but independent of the volume of the space of all tasks. Empirically, we perform a thorough evaluation in high-dimensional environments. We find that the directed goal selection of DISCOVER solves exploration problems that are beyond the reach of prior state-of-the-art exploration methods in RL.
@@ 1 Introduction
Reinforcement learning (RL) provides a general framework in which agents have to learn complex tasks by interacting with their environment to maximize rewards [76]. Although RL has led to breakthroughs in many domains such as Atari games [45], board games [e.g., 66], and reasoning [26], sparse-reward problems, in which a reward is only observed once the task is completed, remain very challenging [38, 81]. These sparse-reward tasks are very hard to solve as the agent needs to gain a deep understanding of the environment without ever observing a reward. As such, and since they are naturally defined without any human supervision, solving sparse-reward tasks is one of the core premises of RL. We focus on the setting where the agent is given a single sparse-reward problem to solve. In sparse-reward RL, traditional methods from RL fail as they need to randomly observe the reward signal to make progress, which becomes exceedingly unlikely in long-horizon tasks [e.g., 49, 88].
@@ 2 Related Work
Solving many sparse-reward problems with RL has often been formalized as goal-conditioned [22, 58] or goal-reaching [84] RL, where we jointly learn a policy for all tasks. A key contribution can be traced back to Schaul et al. [63], which introduces universal value function estimators, capable of estimating returns for multiple goals. To learn these functions efficiently, Andrychowicz et al. [5] proposes a goal relabeling scheme to provide signal for goals that are achieved in hindsight. Further works have explored emerging properties of goal-conditioned value functions, introducing contrastive [22, 9] or quasimetric [84, 18, 3] approaches. Interestingly, goal-conditioned approaches enable the design of nuanced exploration strategies, as the policy can be controlled through its goal-conditioning [56, 86, 53, 23, 70, 87, 35, 43, 55]. These methods strategically select a goal from the previously achieved goals. After the goal is achieved, the methods enter a pure exploration phase, in line with the successful exploration strategy proposed by Ecoffet et al. [20]. However, previous methods are generally not directed towards a particular target goal, with some exceptions relying on strong assumptions on the distance metric [57], or on additional costly components such as generative [40, 62] or discriminative [12, 11] models. In contrast, DISCOVER introduces a notion of relevance during exploration, which is fully bootstrapped from the agent’s own estimates and experience, leading to deep exploration. We further discuss the connection of DISCOVER to other work such as self-play in Appendix A.
@@
We consider the setting where the agent’s objective is to solve a single, challenging target task provided at “test-time”, potentially starting from a pre-trained prior. This is a form of test-time training (TTT) [74], where an agent is trained specifically for the target task at test-time. TTT on (self-)supervised signals over few gradient updates has shown success in domains such as control [32], language modeling [33, 36, 75, 8], abstract reasoning [4], and video generation [15]. More recently, test-time reinforcement learning (TTRL) [89, 68] has used RL to iteratively self-improve an agent’s policy for initially unsolvable tasks, using its own experience on simpler tasks. To our knowledge, DISCOVER is the first method demonstrating effective TTRL with extensive self-supervised exploration (millions of steps), thereby enabling agents to solve highly difficult tasks.
@@ 3.1 Goal Selection for Sparse-Reward RL
3:     $g_{t}\leftarrow\texttt{SelectGoal}(\mathcal{B}_{t-1},\theta_{t-1},g^{\star})$triangleright select goal $g_{t}$ for episode
4:     $\mathcal{B}_{t}\leftarrow\mathcal{B}_{t-1}\cup\{\texttt{Rollout}(\pi_{\theta_{t-1}},g_{t})\}$triangleright add trajectory to replay buffer
5:     $\theta_{t}\leftarrow\texttt{Update}(\texttt{RelabelGoals}(\texttt{Sample}(\mathcal{B}_{t})),\theta_{t-1})$triangleright using off-policy RL (e.g., TD3 [24])
@@ 4 DISCOVER
[…]
Each of these principles is necessary to efficiently learn to solve a hard task. First, achievability ensures that a task is not “too hard” for the agent to ever achieve it, hence, representing meaningful experience. Second, novelty ensures that a task is not “too easy”, such that the agent’s new experience on this task is a useful learning signal to increase the policy’s capabilities. Finally, relevance ensures that experience on a task is useful for eventually solving the target task (cf. Figure 2). Prior methods [such as 53] have successfully combined achievability and novelty, aiming to achieve all possible tasks.
@@
where $\alpha_{t}$ and $\beta_{t}$ are schedulable coefficients, $V$ is the mean of an ensemble of the optimal value function defined in Equation 1, and $\sigma^{2}$ its variance. [^1] See Section E.1. A high value $V(s_{0},g)$ indicates that the policy can reach the goal $g$ from the starting state $s_{0}$, therefore promoting achievability. In contrast, a large uncertainty $\sigma(s_{0},g)$ indicates that the agent is not reliably reaching the goal $g$. Hence, attempting such goals prioritizes novel experiences. Finally, a high value $V(g,g^{\star})$ indicates that $g$ is “closely related” to the target goal $g^{\star}$. The relevance term of DISCOVER prioritizes goals that might be closely related to the target goal, either since they already have a high value $V(g,g^{\star})$ or because the agent is still uncertain about this value. The relevance term can also be viewed as directing the goal selection towards the final target $g^{\star}$. Note that while we focus on a single target $g^{\star}$, DISCOVER is easily extended to distributions over targets by maximizing Equation 3 in expectation.
We emphasize that all terms in Equation 3 are estimated by the critic and do not rely on any prior information. In particular, the relevance estimates are entirely bootstrapped from the agent’s current knowledge. While the above gives an intuitive introduction, DISCOVER can be interpreted as an agent seeking to maximize the likelihood of reaching the target goal $g^{\star}$, as we describe in Appendix B.
@@ 4.1 Automatic Online Parameter Adaptation
Here, $p_{t}$ denotes the average goal achievement rate over the last $k_{\text{adapt}}$ episodes, $\smash{\Pi_{[a,b]}}$ clips the output value to the interval $[a,b]$, and $\eta>0$ is the adaptation step size, which we set to $0.01$. We ablate the effects of our adaptation strategy in Figure 11 in the appendix, and find that the optimal target goal achievement rate $p^{\star}$ is approximately $50\%$, consistent with findings in prior work [53].
@@ 4.2 Connection to the Exploration-Exploitation Dilemma
Informal Assumption 4.1 (formalized in Assumptions C.1, C.2, C.3, C.4 and C.5).
@@
    Goal space contains optimal paths: For any goal $g$, the optimal path from $g$ to the target goal $g^{\star}$ is contained in the goal space $\mathcal{G}$ [^3] That is, $\mathcal{G}$ is $g^{\star}$-geodesically convex under the forward quasimetric induced by the optimal value function.
@@
Assumptions 1 and 2 are standard in the literature on linear bandits. Note, however, that in our setting this feedback model implicitly assumes feedback on $V^{\star}(g_{t},g^{\star})$, which can be thought of as a generalization condition: If the agent cannot generalize from its experience from $s_{0}$ to $g$ about the relation of $g$ and $g^{\star}$, then it may never reach the target goal $g^{\star}$. Another implicit consequence of the linearity assumption is that the bias and non-stationarity of the value function estimates is controlled, which is not commonly the case when learning value functions with neural networks via bootstrapping. Assumption 3 is a continuity assumption on the goal space, which is satisfied in many practical settings. Finally, Assumption 4 leads to a trade-off in choosing $\alpha$: A small $\alpha$ leads to a larger set of achievable goals, but at a lower goal-achieval rate. In contrast, a large $\alpha$ leads to a more conservative set of achievable goals, but at a higher goal-achieval rate. This trade-off matches the empirical effect of $\alpha$ (cf. Figure 11 in the appendix). The parameter $\kappa$ controls the rate of expansion of the set of achievable goals.
@@
Informal Theorem 4.2 (formalized in Theorem C.9).
@@
This theorem shows that DISCOVER efficiently learns to solve hard, initially unachievable tasks. With a larger dimensionality $d$ of the feature space, the learning task becomes harder, which is reflected in the bound. Similarly, the larger the distance $D$ from the initial state to the target goal, the longer it takes to reach the target goal. In contrast, the larger the rate of expansion $\kappa$, the faster the agent reaches the target goal, since the set of achievable goals expands faster. Finally, the bound suggests choosing $\alpha\approx 0.1$ (cf. Figure 8 in the appendix) which roughly matches the average $\alpha$ chosen in experiments by our adaptation strategy (cf. Figure 12 in the appendix).
Despite our simplifying assumptions, achieving the target goal $g^{\star}$ remains non-trivial. The agent must trade off learning the value function against exploiting the direction this value suggests toward $g^{\star}$. Overconfidence in the value estimate may lead the agent to diverge and never reach $g^{\star}$. On the other hand, if the agent is too conservative in its value estimate (or does not use the value estimate for direction at all), it may never reach the target goal within a reasonable time. To see why guidance matters, imagine an $m$-dimensional goal space that is a ball of radius $R$. Covering that ball with hypercubes of side $\epsilon$ requires on the order of $(R/\epsilon)^{m}$ cells, exponential in $m$. This back-of-the-envelope calculation indicates that to achieve hard goals in a high-dimensional goal space, undirected exploration is insufficient, and the agent must balance exploration and exploitation. In doing so, DISCOVER avoids this curse of dimensionality and exploits the learned value to stay close to a nearly one-dimensional corridor from $s_{0}$ to $g^{\star}$. For this reason, the bound in Informal Theorem 4.2 depends solely on the (1-dim) distance $D$ and not on the total volume of the goal space $\mathcal{G}$. Bounds of previous work [79] depend on the total volume of $\mathcal{G}$, and quickly become vacuous in high-dimensional problems.
@@ 5 Results
We evaluate the empirical performance of DISCOVER across three complex, sparse-reward, long-horizon control tasks, highlighting five main insights. Unless mentioned otherwise, we employ the TD3 [24] actor-critic algorithm for training all agents. While our evaluation focuses on model-free methods, the goal-based directed exploration of DISCOVER can also be used with model-based backbones such as Dreamer [29, 30, 31]. For all experiments, we report the mean performance across $10$ seeds along with its standard error. Additional implementation details, hyperparameter choices and experimental results are reported in Appendices E, F and D, respectively. The code is available at https://github.com/LeanderDiazBone/discover.
@@
As shown in Figure 3, DISCOVER consistently outperforms all baseline goal selection strategies across the evaluated environments. The performance gains are particularly notable in the more challenging variants of each task, where emphasizing relevance proves critical for learning optimal policies. Interestingly, in the arm environment, HER performs better than some undirected goal selection strategies. We attribute this to the unconstrained nature of the goal space in the arm environment, which leads undirected methods to explore irrelevant regions that do not contribute to learning about the target goal.
Figure 4: Comparison of the required number of steps for reaching $10\%$ target goal achievement in pointmazes of varying dimension. $\infty$ denotes no achievement of $10\%$ success rate before termination after $50$M steps. DISCOVER scales to large mazes due to its awareness of direction and uncertainty.
@@
We next evaluate the performance of goal selection strategies on pointmaze navigation tasks of varying dimensionality. Figure 4 demonstrates that the undirected goal selection strategies are eventually successful in two dimensions, but consistently fail in dimension larger than three. In contrast, DISCOVER, by focusing on the most relevant directions, successfully solves mazes in up to six dimensions. The substantial difference in empirical performance can be explained by the size of the search space explored by each method. For undirected methods, this space grows exponentially with dimension, while DISCOVER mitigates this challenge by only selecting goals that likely to lead to the final objective, thereby dramatically reducing the effective search space. While the performance gap is most pronounced in complex, high-dimensional tasks, we demonstrate that DISCOVER also improves performance in standard sparse-reward environments.
@@
We visualize the goal selection behavior of DISCOVER and the baseline strategies in the antmaze environment in Figure 5. While all baseline methods perform undirected goal selection, exploring the entire state space, DISCOVER, after an initial exploration phase, quickly identifies the correct direction and subsequently focuses its goal selection on the relevant region. These plots explain the improved performance, as DISCOVER can select (and achieve) goals at the target much earlier than the other strategies. We provide a detailed evaluation of how the different components of DISCOVER influence goal selection in Appendix D (cf. Figure 15).
@@
In many scenarios, prior information about the environment is available, for example through pre-training on a related environment. To evaluate whether DISCOVER can leverage prior knowledge for exploration, we integrate prior information on relevance by substituting the prior for $V(g,g^{\star})$ in Equation 3. We evaluate two kinds of priors: (1) a hand-designed prior, leveraging human knowledge akin to reward shaping, and (2) a pre-trained prior from a similar environment. We use the antmaze (hard) environment, picking as a natural hand-designed prior the $L_{2}$-distance to the target goal $g^{\star}$ (ignoring obstacles), and as the pre-trained prior a value function learned on a pointmaze with the same layout. We find in Figure 6 that DISCOVER with a prior can explore marginally faster than bootstrapping value estimates from scratch. In particular, in the maze with obstacles, using the pre-trained prior accelerates exploration, since the agent explores only in the direction of the target (cf. Figure 14 in the appendix). We expect that this benefit of priors increases in even harder tasks, since DISCOVER explores only the difficult and novel aspects of the target task.
Figure 7: State-of-the-art non-goal-conditioned methods fail to achieve the target tasks even in the “simple” configuration. In contrast, DISCOVER learns to solve these tasks relatively quickly.
@@
Finally, we evaluate the importance of goal-conditioning for solving sparse-reward tasks. To this end, we compare our method to state-of-the-art methods for exploration in RL that do not explicitly select exploratory intermediate goals. This includes approaches based on curiosity and reward shaping [24, 27, 72, 48]. We show in Figure 7 that even in the simpler environments, these standard methods are unable to solve the task. In Figure 16 of the appendix, we visualize the states visited by the non-goal-conditioned baselines. While methods leveraging curiosity to shape the reward explore faster than others, none of them reaches the target task within our maximum number of episodes. This highlights that in long-horizon, sparse-reward settings, directed goal selection can facilitate deep exploration, which is essential for solving complex tasks.

### v2

@@ Abstract
Sparse-reward reinforcement learning (RL) can model a wide range of highly complex tasks. Solving sparse-reward tasks is RL’s core premise—requiring efficient exploration coupled with long-horizon credit assignment—and overcoming these challenges is key for building self-improving agents with superhuman ability. Prior work commonly explores with the objective of solving many sparse-reward tasks, making exploration of individual high-dimensional, long-horizon tasks intractable. We argue that solving such challenging tasks requires solving simpler tasks that are relevant to the target task, i.e., whose achieval will teach the agent skills required for solving the target task. We demonstrate that this sense of direction, necessary for effective exploration, can be extracted from existing RL algorithms, without leveraging any prior information. To this end, we propose a method for directed sparse-reward goal-conditioned very long-horizon RL (DISCOVER), which selects exploratory goals in the direction of the target task. We connect DISCOVER to principled exploration in bandits, formally bounding the time until the target task becomes achievable in terms of the agent’s initial distance to the target, but independent of the volume of the space of all tasks. We then perform a thorough evaluation in high-dimensional environments. We find that the directed goal selection of DISCOVER solves exploration problems that are beyond the reach of prior state-of-the-art exploration methods in RL.
@@ 1 Introduction
Reinforcement learning (RL) provides a general framework in which agents have to learn complex tasks by interacting with their environment to maximize rewards [75]. Although RL has led to breakthroughs in many domains such as Atari games [45], board games [e.g., 66], and reasoning [26], sparse-reward problems, in which a reward is only observed once the task is completed, remain very challenging [38, 80]. These sparse-reward tasks are hard to solve as the agent needs to gain a deep understanding of the environment without ever observing a reward. As such, and since they are naturally defined without any human supervision, solving sparse-reward tasks is one of the core premises of RL. We focus on the setting where the agent is given a single sparse-reward problem to solve. In sparse-reward RL, traditional methods from RL fail as they need to randomly observe the reward signal to make progress, which becomes exceedingly unlikely in long-horizon tasks [e.g., 49, 87].
@@ 2 Related Work
Solving many sparse-reward problems with RL has often been formalized as goal-conditioned [22, 58] or goal-reaching [83] RL, where we jointly learn a policy for all tasks. A key contribution can be traced back to Schaul et al. [63], which introduces universal value function estimators, capable of estimating returns for multiple goals. To learn these functions efficiently, Andrychowicz et al. [5] proposes a goal relabeling scheme to provide signal for goals that are achieved in hindsight. Further works have explored emerging properties of goal-conditioned value functions, introducing contrastive [22, 9] or quasimetric [83, 18, 3] approaches. Interestingly, goal-conditioned approaches enable the design of nuanced exploration strategies, as the policy can be controlled through its goal-conditioning [56, 85, 53, 23, 69, 86, 35, 43, 55]. These methods strategically select a goal from the previously achieved goals. After the goal is achieved, the methods enter a pure exploration phase, in line with the successful exploration strategy proposed by Ecoffet et al. [20]. However, previous methods are generally not directed towards a particular target goal, with some exceptions relying on strong assumptions on the distance metric [57], or on additional costly components such as generative [40, 62] or discriminative [12, 11] models. In contrast, DISCOVER introduces a notion of relevance during exploration, which is fully bootstrapped from the agent’s own estimates and experience, leading to deep exploration. We further discuss the connection of DISCOVER to other work such as self-play in Appendix˜A.
@@
We consider the setting where the agent’s objective is to solve a single, challenging target task provided at “test-time”, potentially starting from a pre-trained prior. This is a form of test-time training (TTT) [73], where an agent is trained specifically for the target task at test-time. TTT on (self-)supervised signals over few gradient updates has shown success in domains such as control [32], language modeling [33, 36, 74, 8], abstract reasoning [4], and video generation [15]. More recently, test-time reinforcement learning (TTRL) [88] has used RL to iteratively self-improve an agent’s policy for initially unsolvable tasks, using its own experience on simpler tasks. To our knowledge, DISCOVER is the first method demonstrating effective TTRL with extensive self-supervised exploration (millions of steps), thereby enabling agents to solve highly difficult tasks.
@@ 3.1 Goal Selection for Sparse-Reward RL
3:  $g_{t}\leftarrow\texttt{SelectGoal}(\mathcal{B}_{t-1},\theta_{t-1},g^{\star})$ $\triangleright$ select goal $g_{t}$ for episode
4:  $\mathcal{B}_{t}\leftarrow\mathcal{B}_{t-1}\cup\{\texttt{Rollout}(\pi_{\theta_{t-1}},g_{t})\}$ $\triangleright$ add trajectory to replay buffer
5:  $\theta_{t}\leftarrow\texttt{Update}(\texttt{RelabelGoals}(\texttt{Sample}(\mathcal{B}_{t})),\theta_{t-1})$ $\triangleright$ using off-policy RL (e.g., TD3 [24])
@@ 4 DISCOVER
$\textbf{{Goal utility}}=\textbf{{Achievability}}+\textbf{{Novelty}}+\textbf{{Relevance}}$ (2)
Each of these principles is necessary to efficiently learn to solve a hard task. First, achievability ensures that a task is not “too hard” for the agent to ever achieve it, hence, representing meaningful experience. Second, novelty ensures that a task is not “too easy”, such that the agent’s new experience on this task is a useful learning signal to increase the policy’s capabilities. Finally, relevance ensures that experience on a task is useful for eventually solving the target task (cf. Figure˜2). Prior methods [e.g., 53] have successfully combined achievability and novelty, but are aiming to achieve all possible tasks.
@@
where $\alpha_{t}$ and $\beta_{t}$ are schedulable coefficients, $V$ is the mean of an ensemble of the optimal value function defined in Equation 1, and $\sigma^{2}$ its variance. [^1] See Section E.1. A high value $V(s_{0},g)$ indicates that the policy can reach the goal $g$ from the starting state $s_{0}$, therefore promoting achievability. In contrast, a large uncertainty $\sigma(s_{0},g)$ indicates that the agent is not reliably reaching the goal $g$. Hence, attempting such goals prioritizes novel experiences. Finally, a high value $V(g,g^{\star})$ indicates that $g$ is “closely related” to the target goal $g^{\star}$. The relevance term of DISCOVER prioritizes goals that might be closely related to the target goal, either since they already have a high value $V(g,g^{\star})$ or because the agent is still uncertain about this value. The relevance term can also be viewed as directing the goal selection towards the final target $g^{\star}$. Note that while we focus on a single target $g^{\star}$, DISCOVER is easily extended to distributions over targets by maximizing Equation˜3 in expectation.
We emphasize that all terms in Equation˜3 are estimated by the critic and do not rely on any prior information. In particular, the relevance estimates are entirely bootstrapped from the agent’s current knowledge. While the above gives an intuitive introduction, DISCOVER can be interpreted as an agent seeking to maximize the likelihood of reaching the target goal $g^{\star}$, as we describe in Appendix B.
@@ 4.1 Automatic Online Parameter Adaptation
Here, $p_{t}$ denotes the average goal achievement rate over the last $k_{\text{adapt}}$ episodes, $\smash{\Pi_{[a,b]}}$ clips the output value to the interval $[a,b]$, and $\eta>0$ is the adaptation step size, which we set to $0.01$. We ablate the effects of our adaptation strategy in Figure˜10 in the appendix, and find that the optimal target goal achievement rate $p^{\star}$ is approximately $50\%$, consistent with findings in prior work [53].
@@ 4.2 Connection to the Exploration-Exploitation Dilemma
Informal Assumption 4.1 (formalized in Assumptions˜C.1, C.2, C.3, C.4 and C.5).
@@
    Goal space contains optimal paths: For any goal $g$, the optimal path from $g$ to the target goal $g^{\star}$ is contained in the goal space $\mathcal{G}$. [^3] That is, $\mathcal{G}$ is $g^{\star}$-geodesically convex under the forward quasimetric induced by the optimal value function.
@@
Assumptions 1 and 2 are standard in the literature on linear bandits. Note, however, that in our setting this feedback model implicitly assumes feedback on $V^{\star}(g_{t},g^{\star})$, which can be thought of as a generalization condition: If the agent cannot generalize from its experience from $s_{0}$ to $g$ about the relation of $g$ and $g^{\star}$, then it may never reach the target goal $g^{\star}$. Another implicit consequence of the linearity assumption is that the bias and non-stationarity of the value function estimates is controlled, which is not commonly the case when learning value functions with neural networks via bootstrapping. Assumption 3 is a continuity assumption on the goal space, which is satisfied in many practical settings. Finally, Assumption 4 leads to a trade-off in choosing $\alpha$: A small $\alpha$ leads to a larger set of achievable goals, but at a lower goal-achieval rate. In contrast, a large $\alpha$ leads to a more conservative set of achievable goals, but at a higher goal-achieval rate. This trade-off matches the empirical effect of $\alpha$ (cf. Figure˜10 in the appendix). The parameter $\kappa$ controls the rate of expansion of the set of achievable goals.
@@
Informal Theorem 4.2 (formalized in Theorem˜C.9).
@@
This theorem shows that DISCOVER efficiently learns to solve hard, initially unachievable tasks. With a larger dimensionality $d$ of the feature space, the learning task becomes harder, which is reflected in the bound. Similarly, the larger the distance $D$ from the initial state to the target goal, the longer it takes to reach the target goal. In contrast, the larger the rate of expansion $\kappa$, the faster the agent reaches the target goal, since the set of achievable goals expands faster. Finally, the bound suggests choosing $\alpha\approx 0.1$ (cf. Figure˜7 in the appendix) which roughly matches the average $\alpha$ chosen in experiments by our adaptation strategy (cf. Figure˜11 in the appendix).
Despite our simplifying assumptions, achieving the target goal $g^{\star}$ remains non-trivial. The agent must trade off learning the value function against exploiting the direction this value suggests toward $g^{\star}$. Overconfidence in the value estimate may lead the agent to diverge and never reach $g^{\star}$. On the other hand, if the agent is too conservative in its value estimate (or does not use the value estimate for direction at all), it may never reach the target goal within a reasonable time. To see why guidance matters, imagine an $m$-dimensional goal space that is a ball of radius $R$. Covering that ball with hypercubes of side $\epsilon$ requires on the order of $(R/\epsilon)^{m}$ cells, exponential in $m$. This back-of-the-envelope calculation indicates that to achieve hard goals in a high-dimensional goal space, undirected exploration is insufficient, and the agent must balance exploration and exploitation. In doing so, DISCOVER avoids this curse of dimensionality and exploits the learned value to stay close to a nearly one-dimensional corridor from $s_{0}$ to $g^{\star}$. For this reason, the bound in Informal Theorem˜4.2 depends solely on the (1-dim) distance $D$ and not on the total volume of the goal space $\mathcal{G}$. Bounds of previous work [78] depend on the total volume of $\mathcal{G}$, and quickly become vacuous in high-dimensional problems.
@@ 5 Results
We evaluate the empirical performance of DISCOVER across three complex, sparse-reward, long-horizon control tasks, highlighting five main insights. Unless mentioned otherwise, we employ the TD3 [24] actor-critic algorithm for training all agents. While our evaluation focuses on model-free methods, the goal-based directed exploration of DISCOVER can also be used with model-based backbones such as Dreamer [29, 30, 31]. For all experiments, we report the mean performance across $10$ seeds along with its standard error. Additional implementation details, hyperparameter choices and experimental results are reported in Appendices˜E, E.3 and D, respectively. The code is available at https://github.com/LeanderDiazBone/discover.
@@
As shown in Figure˜3, DISCOVER consistently outperforms all baseline goal selection strategies across the evaluated environments. The performance gains are particularly notable in the more challenging variants of each task, where emphasizing relevance proves critical for learning optimal policies. Interestingly, in the arm environment, HER performs better than some undirected goal selection strategies. We attribute this to the unconstrained nature of the goal space in the arm environment, which leads undirected methods to explore irrelevant regions that do not contribute to learning about the target goal.
Dimension 2 3 4 5 6
HER $\infty$ $\infty$ $\infty$ $\infty$ $\infty$
MEGA 4.8 $\infty$ $\infty$ $\infty$ $\infty$
Ach. + Nov. 5.2 $\infty$ $\infty$ $\infty$ $\infty$
[…]
Table 1: Comparison of the required number of steps (M) for reaching $10\%$ target goal achievement in pointmazes of varying dimension. $\infty$ denotes no achievement of $10\%$ success rate before termination after 5050M steps. DISCOVER scales to large mazes due to its awareness of direction and uncertainty.
@@
We next evaluate the performance of goal selection strategies on pointmaze navigation tasks of varying dimensionality. Table˜1 demonstrates that the undirected goal selection strategies are eventually successful in two dimensions, but consistently fail in dimension larger than three. In contrast, DISCOVER, by focusing on the most relevant directions, successfully solves mazes in up to six dimensions. The substantial difference in empirical performance can be explained by the size of the search space explored by each method. For undirected methods, this space grows exponentially with dimension, while DISCOVER mitigates this challenge by only selecting goals that likely to lead to the final objective, thereby dramatically reducing the effective search space. While the performance gap is most pronounced in complex, high-dimensional tasks, we demonstrate that DISCOVER also improves performance in standard sparse-reward environments.
@@
We visualize the goal selection behavior of DISCOVER and the baseline strategies in the antmaze environment in Figure˜4. While all baseline methods perform undirected goal selection, exploring the entire state space, DISCOVER, after an initial exploration phase, quickly identifies the correct direction and subsequently focuses its goal selection on the relevant region. These plots explain the improved performance, as DISCOVER can select (and achieve) goals at the target much earlier than the other strategies. We provide a detailed evaluation of how the different components of DISCOVER influence goal selection in Appendix˜D (cf. Figure˜14).
@@
In many scenarios, prior information about the environment is available, for example through pre-training on a related environment. To evaluate whether DISCOVER can leverage prior knowledge for exploration, we integrate prior information on relevance by substituting the prior for $V(g,g^{\star})$ in Equation˜3. We evaluate two kinds of priors: (1) a hand-designed prior, leveraging human knowledge akin to reward shaping, and (2) a pre-trained prior from a similar environment. We use the antmaze (hard) environment, picking as a natural hand-designed prior the $L_{2}$-distance to the target goal $g^{\star}$ (ignoring obstacles), and as the pre-trained prior a value function learned on a pointmaze with the same layout. We find in Figure˜5 that DISCOVER with a prior can explore marginally faster than bootstrapping value estimates from scratch. In particular, in the maze with obstacles, using the pre-trained prior accelerates exploration, since the agent explores only in the direction of the target (cf. Figure˜13 in the appendix). We expect that this benefit of priors increases in even harder tasks, since DISCOVER explores only the difficult and novel aspects of the target task.
Figure 6: State-of-the-art RL algorithms fail to achieve the target tasks even in their “simple” configuration.
@@
Finally, we evaluate the importance of goal-conditioning for solving sparse-reward tasks. To this end, we compare DISCOVER to methods for exploration in RL that do not explicitly select exploratory intermediate goals [24, 27, 48]. This includes state-of-the-art approaches based on curiosity and reward shaping [10, 71]. We show in Figure˜6 that even in the simpler environments, these standard methods are unable to solve the task. In Figure˜15 of the appendix, we visualize the states visited by the non-goal-conditioned baselines. While methods leveraging curiosity to shape the reward explore faster than others, none of them reaches the target task within our maximum number of episodes. This highlights that in long-horizon, sparse-reward settings, directed goal selection can facilitate deep exploration, which is essential for solving complex tasks.